### v1

# Unifying Decoherence and Phase Evolution in Mixed Quantum–Classical Dynamics through Exact Factorization

Abstract
We propose new mixed quantum–classical equations of motion within the exact factorization framework to describe electronic coherence and phase evolution simultaneously. The derivation reveals that not only a projected quantum momentum correction, recently identified by Arribas and Maitra [Phys. Rev. Lett. 133, 233201 (2024)], but also a phase-correction appears within the order of $\hbar$, providing a unified and rigorous account of electronic coherence and phase evolution as well as their effect on the nuclear force. Benchmark tests on one- and two-dimensional model systems demonstrate that the new formulations capture key nonadiabatic features with high accuracy.
Understanding the coupled motion of electrons and nuclei in molecules remains one of the central challenges in theoretical chemistry and chemical physics. Fully quantum treatments of electron–nuclear dynamics are computationally prohibitive for systems of realistic size, owing to the scaling of the many-body molecular wavefunction. To overcome this limitation, mixed quantum–classical (MQC) methods [1, 2, 3] where the electrons are treated quantum mechanically while nuclei follow classical trajectories have become indispensable tools for simulating photochemical processes, charge and energy transfer, and ultrafast dynamics in molecules and materials [4, 5, 6, 7, 8, 9]. Among MQC methods, Ehrenfest dynamics [1] and trajectory-based surface hopping [2] are the most widely applied. Despite their utility, they suffer from two fundamental shortcomings: the absence of a reliable mechanism for electronic decoherence, and the lack of a rigorous treatment of phase evolution between adiabatic states. Numerous correction schemes have been proposed, often guided by semiclassical arguments, empirical adjustments, intuition from wave-packet models, or mapping of variables [10, 11, 12, 13, 14, 15, 16, 17].
The exact factorization (XF) formalism provides a rigorous foundation for electron–nuclear dynamics by expressing the molecular wavefunction as a product of a nuclear wavefunction and a conditional electronic state [18, 19]. This leads to coupled electronic and nuclear equations of motion that, in principle, include the full electron–nuclear correlation (ENC) effects. XF-based approaches have therefore attracted attention as a systematic framework for deriving improved MQC dynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In recent work, Arribas and Maitra showed that a projected quantum momentum (PQM) correction naturally arises from the XF formalism and plays a central role in describing electronic coherence [32].
In this work, we derive new mixed quantum–classical equations of motion from the exact factorization formalism, revealing an additional phase-correction term that accompanies the PQM correction of comparable or smaller order in $\hbar$. This previously unidentified term governs the proper phase evolution of the Born–Oppenheimer (BO) coefficients. The combined inclusion of PQM and phase corrections provides a balanced and rigorous description of both decoherence and phase dynamics without relying on semiclassical approximations or heuristic arguments, offering deeper conceptual insight into ENC and a practical route to accurate molecular quantum dynamics. We validate the equations through benchmark calculations for the double-arch geometry (DAG) [14] and two-dimensional nonseparable (2DNS) [15] models, which demonstrate that both corrections are essential for reproducing characteristic nonadiabatic features such as Stückelberg oscillations, intermediate-time electronic coherence, and the correct nuclear-density distributions.
The XF formalism expresses the molecular wavefunction $|\Psi({\underline{\underline{\textbf{R}}}},t)\rangle$ as a single product of a time-dependent nuclear wavefunction $\chi({\underline{\underline{\textbf{R}}}},t)$ and an electronic wavefunction $|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle$, $|\Psi({\underline{\underline{\textbf{R}}}},t)\rangle=\chi({\underline{\underline{\textbf{R}}}},t)|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle$ with partial normalization condition $\langle\Phi_{\underline{\underline{\textbf{R}}}}(t)|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle_{e}=1$ at any $\underline{\underline{\textbf{R}}}$} and $t$, where $\underline{\underline{\textbf{R}}}$} represents the nuclear degrees of freedom, and $\langle\cdots\rangle_{e}$ represents an integral over electronic degrees of freedom [18, 19]. Then, from the coupled time-dependent Schrödinger equations (TDSEs) for $\chi({\underline{\underline{\textbf{R}}}},t)$ and $|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle$, the MQC equations can be derived [21, 29]. The exact classical force for the nuclear trajectory ${\underline{\underline{\textbf{R}}}}(t)$ is
$\begin{split}\mathbf{F}_{\nu}=\dot{\mathbf{P}}_{\nu}(t)=-\nabla_{\nu}\tilde{\epsilon}+\dot{\textbf{A}}_{\nu},\end{split}$ (1)
and the electronic TDSE along the nuclear trajectory is
$i\hbar\dfrac{d}{dt}|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle=\left(\hat{H}_{\mathrm{BO}}+\hat{H}^{(1)}_{\mathrm{ENC}}+\hat{H}^{(2)}_{\mathrm{ENC}}-\tilde{\epsilon}\right)|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle,$ (2)
where $\hat{H}_{\mathrm{BO}}$ is the BO Hamiltonian, and $\textbf{A}_{\nu}({\underline{\underline{\textbf{R}}}},t)=\langle\Phi_{\underline{\underline{\textbf{R}}}}|-i\hbar\nabla_{\nu}\Phi_{\underline{\underline{\textbf{R}}}}\rangle_{e}$ and $\tilde{\epsilon}({\underline{\underline{\textbf{R}}}},t)=\langle\Phi_{\underline{\underline{\textbf{R}}}}(t)|(\hat{H}_{\mathrm{BO}}+\hat{H}^{(2)}_{\mathrm{ENC}}-i\hbar\tfrac{d}{dt})|\Phi_{\underline{\underline{\textbf{R}}}}(t)\rangle_{e}$ are the time-dependent vector and scalar potentials, respectively. Here, $\dot{f}$ represents the total time-derivative $df/dt$. The first and second electron–nuclear correlation (ENC) operators are $\hat{H}^{(1)}_{\mathrm{ENC}}=-\sum_{\nu}\dfrac{\hbar^{2}}{2M_{\nu}}\bm{\mathcal{P}}_{\nu}\cdot\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}$ and $\hat{H}^{(2)}_{\mathrm{ENC}}=-\sum_{\nu}\dfrac{\hbar^{2}}{2M_{\nu}}\left(\nabla_{\nu}^{2}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}+\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}\cdot\nabla_{\nu}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}\right)$ where $\bm{\mathcal{P}}_{\nu}=\nabla_{\nu}|\chi|^{2}/|\chi|^{2}$ is the total nuclear quantum momentum (QM), and $\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}=|\Phi_{\underline{\underline{\textbf{R}}}}\rangle\langle\Phi_{\underline{\underline{\textbf{R}}}}|$ is the reduced electron density operator at a nuclear configuration $\underline{\underline{\textbf{R}}}$}. We note that all terms are evaluated along ${\underline{\underline{\textbf{R}}}}(t)$.
Further expansion of Eq. (2) in terms of BO basis states $|\Phi_{{\underline{\underline{\textbf{R}}}}}(t)\rangle=\sum_{j}C_{j}|\phi_{j}\rangle=\sum_{j}|C_{j}|e^{iS_{j}/\hbar}|\phi_{j}\rangle$ satisfying $\hat{H}_{\mathrm{BO}}|\phi_{j}\rangle=\epsilon_{j}|\phi_{j}\rangle$ yields
$\dot{C}_{j}=\dot{C}_{j}^{\mathrm{Eh}}+\dot{C}_{j}^{\mathrm{QM}}+\dot{C}_{j}^{\mathrm{PQM}}+\dot{C}_{j}^{\mathrm{Div}}+\dot{C}_{j}^{\mathrm{Ph}}.$ (3)
For the time evolution of BO coefficients, we have contributions from the Ehrenfest term $\dot{C}_{j}^{\mathrm{Eh}}=-\frac{i}{\hbar}\epsilon_{j}C_{j}-\sum_{\nu,k}\frac{\textbf{P}_{\nu}}{M_{\nu}}\cdot\textbf{d}_{\nu,jk}C_{k}$, the QM term $\dot{C}_{j}^{\mathrm{QM}}=-\sum_{\nu,k}\frac{\bm{\mathcal{P}}_{\nu}}{2M_{\nu}}\cdot\mathbf{D}_{\nu,jk}|C_{k}|^{2}C_{j}$ with $\mathbf{D}_{\nu,jk}=\nabla_{\nu}S_{j}-\nabla_{\nu}S_{k}$, the PQM [32] term $\dot{C}_{j}^{\mathrm{PQM}}=-\sum_{\nu,k}\frac{\bm{\mathcal{Q}}_{\nu,jk}}{2M_{\nu}}\cdot\mathbf{D}_{\nu,jk}|C_{k}|^{2}C_{j}$ where $\bm{\mathcal{Q}}_{\nu,jk}=\nabla_{\nu}|C_{j}|^{2}/|C_{j}|^{2}+\nabla_{\nu}|C_{k}|^{2}/|C_{k}|^{2}$, the divergence term $\dot{C}_{j}^{\mathrm{Div}}=-\sum_{\nu,k}\frac{\nabla_{\nu}\cdot\mathbf{D}_{\nu,jk}}{2M_{\nu}}|C_{k}|^{2}C_{j}$, and the phase correction term $\dot{C}_{j}^{\mathrm{Ph}}=\frac{i}{\hbar}(\tilde{\epsilon}-\sum_{\nu}\frac{|\nabla_{\nu}S_{j}-\textbf{P}_{\nu}|^{2}}{2M_{\nu}})C_{j}$. Here we neglect the $O(\hbar^{2})$ terms and the nonadiabatic coupling (NAC) terms ($\textbf{d}_{\nu,jk}=\langle\phi_{j}|\nabla_{\nu}\phi_{k}\rangle_{e}$ and $g_{\nu,jk}=\langle\phi_{j}|\nabla^{2}_{\nu}\phi_{k}\rangle_{e}$) in ENC terms.
The nuclear force [Eq. (1)] in the BO representation can be expressed as
$\mathbf{F}_{\nu}=\mathbf{F}_{\nu}^{\mathrm{Eh}}+\mathbf{F}_{\nu}^{\mathrm{QM}}+\mathbf{F}_{\nu}^{\mathrm{PQM}}+\mathbf{F}_{\nu}^{\mathrm{Div}}+\mathbf{F}_{\nu}^{\mathrm{Ph}}+\mathbf{F}_{\nu}^{\mathrm{GI}}$ (4)
which directly corresponds to the contributions in Eq. (3) except $\mathbf{F}_{\nu}^{\mathrm{GI}}$ from the gauge-invariant part of the time-dependent scalar potential, $\tilde{\epsilon}_{\mathrm{GI}}=\langle\Phi_{\underline{\underline{\textbf{R}}}}|\hat{H}^{(2)}_{\mathrm{ENC}}|\Phi_{\underline{\underline{\textbf{R}}}}\rangle_{e}$ in Eq. (1). Detailed derivations for Eqs. (3) and (4) are provided in Supplemental Material [33].
The conventional XF-based MQC approaches, coupled-trajectory MQC (CTMQC) [20, 21, 22, 23, 24, 25, 26] or independent-trajectory approaches such as surface hopping dynamics based on XF (SHXF) [27] and Ehrenfest dynamics based on XF (EhXF) [28], only consider $\dot{C}_{j}=\dot{C}^{\mathrm{Eh}}_{j}+\dot{C}^{\mathrm{QM}}_{j}$ and $\mathbf{F}_{\nu}=\mathbf{F}_{\nu}^{\mathrm{Eh}}+\mathbf{F}_{\nu}^{\mathrm{QM}}$ obtained from the first ENC operator $\hat{H}_{\mathrm{ENC}}^{(1)}$ where $\dot{C}^{\mathrm{QM}}_{j}$ and $\mathbf{F}_{\nu}^{\mathrm{QM}}$ provide decoherence with nuclear wave-packet splitting. However, these approaches cannot provide accurate phase evolution [21, 28]. Recently, Arribas and Maitra proposed that an additional correction to the total QM, $\dot{C}^{\mathrm{PQM}}_{j}$, is necessary to consider electronic coherence more accurately [32].
Here we claim that there are more additional terms to be considered at the same or less order as the QM correction, the phase term, $\dot{C}^{\mathrm{Ph}}_{j}$, and the divergence term, $\dot{C}^{\mathrm{Div}}_{j}$ which are crucial to describe correct BO phases and coherence simultaneously. These terms as well as $\dot{C}^{\mathrm{PQM}}_{j}$ are derived from the second ENC operator which has been previously overlooked, especially from $\nabla_{\nu}^{2}\hat{\Gamma}_{\underline{\underline{\textbf{R}}}}$ in $\hat{H}^{(2)}_{\mathrm{ENC}}$ [33].
The phase term, $\dot{C}^{\mathrm{Ph}}_{j}$, is of the same order as $\frac{i}{\hbar}\epsilon_{j}C_{j}$ in $\dot{C}^{\mathrm{Eh}}_{j}$, which governs an additional phase evolution. Here, $\tilde{\epsilon}$ is a global phase, and can therefore be neglected during time evolution. If we approximate $\nabla_{\nu}S_{i}$ by the BO momentum $\textbf{P}_{\nu,i}\equiv\alpha_{i}\textbf{P}_{\nu}$, thus $\mathbf{D}_{\nu,jk}=\textbf{P}_{\nu,j}-\textbf{P}_{\nu,k}$, with $\alpha_{j}$ determined by the state-wise total energy conservation condition, $\alpha_{i}=\sqrt{\max\left(0,\frac{E_{\mathrm{tot}}-\epsilon_{i}}{\sum_{\mu}|\textbf{P}_{\mu,i}|^{2}/M_{\mu}}\right)}$, we obtain
$\begin{split}\dot{S}_{i}-\dot{S}_{j}=&-\sum_{\nu}\dfrac{\textbf{P}_{\nu,j}-\textbf{P}_{\nu,k}}{M_{\nu}}\cdot\textbf{P}_{\nu}\end{split}$ (5)
from phases in $\dot{C}^{\mathrm{Eh}}$ and $\dot{C}^{\mathrm{Ph}}$. This shows that the semiclassical phase correction that had previously been introduced in PCSH heuristically [15] is actually an approximation of the phase term of XF. In this context, we expect that the Stückelberg oscillations can be captured with $\dot{C}^{\mathrm{Ph}}_{j}$ as PCSH.
The divergence contribution, $\dot{C}^{\mathrm{Div}}_{j}$, contains the divergence $\mathbf{D}_{\nu,jk}$, which requires the second derivatives of phases. Despite the difficulty to obtain the second derivatives, the $\dot{C}^{\mathrm{Div}}_{j}$ is essential for the conservation of BO population in the absence of NAC, i.e. $\int d{\underline{\underline{\textbf{R}}}}\dot{|C_{j}|^{2}}|\chi|^{2}=\int\nabla_{\nu}\cdot(\sum_{k}|C_{j}|^{2}|C_{k}|^{2}|\chi|^{2}\mathbf{D}_{\nu,jk})=0$, which is obtained from $\dot{C}^{\mathrm{QM}}+\dot{C}^{\mathrm{PQM}}+\dot{C}^{\mathrm{Div}}$ since $\dot{C}^{\mathrm{Eh}}$ without NAC and $\dot{C}^{\mathrm{Ph}}$ does not contribute to the BO population transfer. Therefore, in this work, we approximate the divergence term using constant pair-wise shift terms (See Supplemental Material [33] for the numerical implementation of the equations).
We benchmark the new equations of motion using the one-dimensional DAG [14] and the 2DNS [15] models. The model Hamiltonians and computational details are provided in the Supplemental Material [33].
Figure 1: Transmission probability for the lower state ($T_{1}$) as a function of initial momentum $k_{0}$ for various dynamics methods. $T_{1}$ is obtained as $T_{1}=\int_{0}^{\infty}dR\;|\chi_{1}(R,t_{f})|^{2}$ for the exact wave-packet dynamics and $T_{1}=\sum_{I\in\{R^{(I)}(t_{f})>0\}}|C_{1}^{(I)}(t_{f})|^{2}/N_{\mathrm{traj}}$ for the trajectory-based dynamics. $R^{(I)}(t)$ and $t_{f}$ denote the trajectory as a function of time $t$ and the final time of a simulation, respectively, with the total number of trajectories $N_{\mathrm{traj}}$.
Fig. 1 shows the transmission probability on the lower state, $T_{1}$, for the DAG model with various trajectory-based approaches in comparison with the exact wave-packet dynamics. We refer to the results obtained without the additional terms as SHXF and CT, whereas those including the term are denoted as SHXFv2 and CTv2, respectively. The exact behavior shows an increase of oscillation amplitude as $k_{0}$ increases, from $\sim$ 0.5 at $k_{0}=30.0$ a.u. As reported previously, PCSH can correctly reproduce the exact Stückelberg oscillation as aimed [15]. On the other hand, PCSH with the energy-based decoherence correction [11] (PCSH-EDC) removes the oscillation due to over-decoherence. Similarly, CT, which only account for the decoherence, shows $T_{1}\sim 0.5$ for all values of $k_{0}$. Although SHXF shows an oscillatory behavior with respect to $k_{0}$, it deviates from the exact behavior. CTv2 and SHXFv2 recover the correct Stückelberg oscillation. While CTv2 slightly underestimates the overall amplitude, SHXFv2 shows excellent agreement with the exact result at high $k_{0}$, and is qualitatively correct at low $k_{0}$. This demonstrates that the additional phase correction term is essential for the correct time evolution of BO phases.
Figure 2: Time evolution of BO population $\langle|C_{1}|^{2}\rangle(t)$ (upper) and coherence $\langle|C_{1}C_{2}|^{2}\rangle(t)$ (lower) with different initial momenta $k_{0}=30$ and $40$ a.u.
Although reproducing the correct oscillation in $T_{1}$ is encouraging, it remains essential to evaluate the time propagation of wave packets to fully assess the accuracy. Fig. 2 shows the time evolution of BO population for the lower state, $\langle|C_{1}|^{2}\rangle(t)$, and coherence, $\langle|C_{1}C_{2}|^{2}\rangle(t)$, with different initial momenta $k_{0}=30.0$ and $50.0$ a.u. For exact wave-packet dynamics, $\langle O\rangle(t)$ is defined as $\langle O\rangle(t)=\int d{\underline{\underline{\textbf{R}}}}\;O({\underline{\underline{\textbf{R}}}},t)$, whereas $\langle O\rangle(t)\equiv\sum_{I}^{N_{\mathrm{traj}}}O^{(I)}(t)/N_{\mathrm{traj}}$ is used for MQC methods. Since the DAG model contains two NAC regions, the initial population drop corresponds to passage through the first region, while the separated wave packets on upper and lower states afterward encounter the second region at different times.
PCSH reproduces the final BO populations well but fails to capture the intermediate dynamics for $k_{0}=30.0$ and shows overcoherence for all $k_{0}$ due to the absence of a decoherence correction. In contrast, PCSH-EDC exhibits over-decoherence, yielding final $\langle|C_{1}|^{2}\rangle$ to $\sim 0.5$ for all $k_{0}$. All XF-based MQC methods reasonably reproduce the correct BO population dynamics for $k_{0}=30.0$ a.u. Notably, only CTv2 and SHXFv2 reproduce the accurate intermediate decoherence at $t\sim 2000$ a.u., highlighting the importance of the additional terms in CTv2 and SHXFv2. For $k_{0}=40.0$ a.u., CT incorrectly predicts the final population to $\sim$0.5, while SHXF yields the opposite direction of second population transfer. Again, SHXFv2 and CTv2 successfully reproduce the correct BO population and coherence dynamics, including the population transfer at the second NAC region. In these cases, CTv2 provides better coherence between the first and second NAC regions, whereas SHXFv2 exhibits slight overcoherence.
Figure 3: Spatial distribution of nuclear wave packets ($|\chi_{i}(R,t)|^{2}$) and BO populations ($|C_{i}(R,t)|^{2}$) at different time steps, $t=1300$ a.u. (a,c,e,g) and $1900$ a.u. (b,d,f,h) with the initial momentum $k_{0}=50.0$ a.u. for CT-based approaches (CT/CTv2) (a-d) and SHXF-based approaches (SHXF/SHXFv2) (e-h). The BO potential energies $\epsilon_{1/2}$ and the NAC $d_{12}$ are depicted to interpret wave-packet dynamics (a,b,e,f).
In addition, to analyze the wave-packet dynamics, we construct the spatial distribution of the BO-projected time-dependent nuclear densities, $|\chi_{i}(R,t)|^{2}$, and BO populations, $|C_{i}(R,t)|^{2}$, for the DAG model. $|\chi_{i}(R,t)|^{2}$ are obtained from histograms of nuclear trajectories weighted by their corresponding BO populations, while $|C_{i}(R,t)|^{2}$ are calculated by $|C_{i}(R,t)|^{2}=|\chi_{i}(R,t)|^{2}/|\chi(R,t)|^{2}$ where the total nuclear density, $|\chi(R)|^{2}$, is the histograms of trajectories. Fig. 3 presents $|\chi_{i}(R,t)|^{2}$ and $|C_{i}(R,t)|^{2}$ from the XF-based approaches for large $k_{0}$ ($k_{0}=50$ a.u.), at $t=1300$ a.u. (arrival of nuclear wave packets at the second NAC region) and $t=1900$ a.u. (after passage through the second NAC region). We observe a marked improvement in the spatial distributions with CTv2 and SHXFv2. CT exhibits oscillatory behavior arising from incorrect phase estimation in each BO state, while SHXF yields an overall inaccurate spatial distribution. By contrast, CTv2 reproduces almost identical results to the exact dynamics. SHXFv2 performs reasonably well at $k=1900$ a.u., but at $k=1300$ a.u. it produces strongly overlapping nuclear wave packets, likely due to a less accurate treatment of decoherence.
Figure 4: Spatial distribution of nuclear wave packets (a,b) and BO populations (c,d) at $t=2200$ a.u. with the initial momentum $k_{0}=30.0$ a.u. for CTv2 (a,c) and CTv2 with the QM $\mathbf{\mathcal{P}}_{\nu}$ and the PQM $\mathbf{\mathcal{Q}}_{\nu}$ obtained from quantum dynamics (b,d). BO potentials and NAC are depicted as in Fig. 3.
For $k_{0}=30.0$ a.u., the nuclear wave-packet splitting and the electronic decoherence effects are pronounced after passing the first NAC region because of the small kinetic energy, due to a larger relative difference in the velocities of nuclear wave packets on different BO states. Thus, the well-separated nuclear wave packets pass the second NAC region sequentially. Fig. 4 shows nuclear wave packets and BO populations at $t=2200$ a.u. when $|\chi_{1}(R,t)|^{2}$ meets the second NAC region. CTv2 provides better wave-packet splitting and $|C_{i}(R,t)|^{2}$ compared to CT in Fig. LABEL:fig:ct_spatial_k0_30 [33], while both deviate from the exact result. On the other hand, the CTv2 with the exact QM and PQM recovers the correct $|\chi_{i}(R,t)|^{2}$ and $|C_{i}(R,t)|^{2}$, whereas CT [Fig. LABEL:fig:ct_spatial_k0_30] with the exact QM still deviates from the exact result [33]. Thus, the formulation is valid; the discrepancies mainly originate from the approximations to $\bm{\mathcal{P}}_{\nu}$ and $\bm{\mathcal{Q}}_{\nu,ik}$, as the true $|\chi_{2}(R,t)|^{2}$ becomes strongly non-Gaussian after the first NAC region for $k_{0}=30.0$ a.u.
The results from other approaches are provided in Figs. LABEL:fig:ct_spatial_k0_30LABEL:fig:pcsh_both_spatial_k0_50 [33]. For all $k_{0}$, PCSH and PCSH-EDC yields either incorrect spatial decoherence or incorrect BO population after the second NAC region, respectively, whereas CTv2 and SHXFv2 exhibit clear enhancement compared to the other methods.
Figure 5: Transmission probability for the lower state ($T_{1}$) for the 2DNS model, as in Fig. 1.
Finally, to test the additional corrections in higher-dimensional systems, we compute the $T_{1}$ for the 2DNS model. As shown in Fig. 5, SHXFv2 and CTv2, along with PCSH, reproduce the correct oscillation pattern, whereas PCSH-EDC and SHXF fail, consistent with the DAG results. CT works for $k_{0}>25$ a.u. but fails for $k_{0}=17$–25 a.u. Thus, the phase and PQM corrections enable us to reproduce the quantum dynamics even in higher dimensions.
In this work, we developed new mixed quantum–classical equations of motion for molecules within the exact factorization formalism. The derivation reveals that additional terms of the same or smaller order in $\hbar$ naturally arise from the full electron–nuclear correlation. In contrast to existing correction schemes, where decoherence and phase adjustments are often introduced heuristically and separately, the exact factorization framework provides both effects simultaneously and without ad hoc assumptions. The newly identified terms consist of the projected quantum momentum, which governs decoherence by correctly separating electronic populations, and a phase-correction term, which ensures accurate phase evolution of the nuclear wave packets.
Numerical benchmarks on representative one- and two-dimensional models demonstrate that these corrections are indispensable for reproducing key features of quantum dynamics, including Stückelberg oscillations, intermediate-time (de)coherence, and the correct spatial distribution of nuclear wave packets. In particular, the enhanced performance of CTv2 and SHXFv2 underscores the importance of explicitly including both decoherence and phase corrections when simulating coupled electron–nuclear motion. Thus, our results establish that the exact factorization formalism provides not only a rigorous foundation for trajectory-based dynamics but also a systematic pathway for deriving corrections that capture the full scope of electron–nuclear correlation effects.
Acknowledgements.
This research was supported by the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT (MSIT)) (RS-2023-NR119931, RS-2023-00257666, RS-2024-00455131).